Lifetimes of a certain brand of lightbulbs is known to follow a right-skewed distribution with mean 24 months and standard deviation 2 months. Suppose that a sample of 14 lightbulbs is taken. What is the probability that the average lifetime of these bulbs is less than 23.5 months?

Select one:

a. Not enough information has been given to answer the question.

b. 0.1736

c. 0.0000

d. 0.0000

e. 0.8264

f. 0.3472
Answer #1

Solution :

Given that ,

mean = = 24

standard deviation = = 2

n = 14

=    = 24

= / n = 2 / 14 = 0.53

P( < 23.5) = P(( - ) / < (23.5 - 24) /0.53 )

= P(z < -0.94)

Using z table

= 0.1736

correct option is = b